Please give detailed steps. Thank you.

3.4 Utility IV (Challenging) Consider an agent whose utility function u(x,X2- + x1x2 + expresses her a) Compute the marginal rate of substitution. b) Draw two or more of the agents indifference curves
Answer #1

u = x12 + x1x2 + x22

(a) MRS = MU1/MU2

MU1 = partial u/partialx1 = 2x1 + x2

MU2 = partial u/partialx2 = x1 + 2x2

MRS = (2x1 + x2) / (x1 + 2x2)

(b) Indifference curves are as follows. The farther away from origin the indifference curve is, the higher the total utility.

IC IC ICI 3

(c) Preference is convex if partial 2u/partialx12 > 0 and partial 2u/partialx22 > 0.

partial2u/partialx12 = partial /partialx1(partialu/partialx1) = partial /partialx1(2x1 + x2) = 2 > 0

partial2u/partialx22 = partial /partialx2(partialu/partialx2) = partial /partialx2(x1 + 2x2) = 2 > 0

Therefore preferences are convex.